OpenStudy (anonymous):

Solve the following inequality. -9/20a-5>4

OpenStudy (anonymous):

First, we must normalize it. \[-\frac{9}{20}a - 5 >4\] We want to solve for the variable, i.e. \[a\]... so let's put it in terms of the variable. Add 5 to both sides. \[-\frac{9}{20}a > 9\] Now, let's divide both sides by \[-\frac{9}{20}\] or, equivalently, multiply by \[-\frac{20}{9}\] So, see as follows. Note that when multiplying by -1, we need to take care of flipping the equivalence relation. \[a < -20\]